find the first three terms and the 100th term of sequences whose nth term is given by:
10n-1

Found 2 solutions by luke94, solver91311:
Answer by luke94(89)

hi there, well the first three terms (starting at 0) and 100th term is
first term: 10*0 - 1 = -1
second term: 10*1 - 1 = 9
third term: 10*2 - 1 = 99
hundreth term: 10*100 - 1 = 999
just substitute whatever term you are finding as your 'n' value

Answer by solver91311(24713)



Same process as the other problems just like this one.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism